1.

3 371\begin{bmatrix} x\\ y\\ z\\ \end{bmatrix}10 22.50 35.50

a) Find the LU factorization of A.

b) Use this factorization to determine the cost of bushel of each type of grain. (Hint: L(Ux) = b.) State your answer in the form "one bushel of high-grade grain costs $x", etc.

c) What is the exact linear combination of the column of A that gives b? What geometric objects (lines in R2, planes in R3, 3-D spaces, etc.) are represented by each equation in the system, and in what ways do these objects intersect? (In other words, describe the column and row pictures for this system Ax =b.)

d) In the market the next day the grain-seller has raised the prices. Now, the first bundle costs $15, the second bundle costs $45, and the third bundle costs $70. Without actually trying to solve this new system, determine whether it will be possible to find a solution x to the new system Ax = (15,45,70). Explain your answer. (Hint: the column space is composed of all right-hand sides b so that Bx = b is solvable.)

3 371

10 22.50 35.50
Answer #1

1 3 210 2 7 5 y 22.50 3 11 9 35.50 Answer: a) Find the LU factorization of A: 1 001 [u11 u12 u13 A-2 7 5 121 1 00 u22 u23 3 1131u12 + 13H42-11, or, or, l3z 3 X3 + 132-11 31 13 or Thus LU factorization of A 1 0 01 1 3 2 3 119 3 2 1J L0 0 1 b) Use thisThus u10,v- 22.50 - 2u-2.50, w = 35.50-3:1-20 = 35.50-30-5 0.50 Then we solve 1 3 21 x 10 1||.yl=!v!= | 2.5 WL0.5 1 Thus z=0.c) The exact linear combination of the column of A that gives b is 10 3 |2| + 2 | 7 | + 0.5 |5-122.50-b 3 11 91 135.50 Each e